Ellen wishes to mix candy worth $1.56 per pound with candy worth $2.60 per pound to form 20 pounds of a mixture worth $2.29 per pound. How many pounds of the more expensive candy should she use?
Answer by checkley71(8403)

2.6x+1.56(20-x)=2.29*20
2.6x+31.2-1.56x=45.8
1.04x=45.8-31.2
1.04x=14.6
x=14.6/1.04
x=14.04 pounds of the expensive candy is required.
PROOF:
2.6*14.04+1.56(20-14.04)=45.8
36.50+1.56*5.96=45.8
36.50+9.30=45.8
45.8=45.8